# Heterogeneous Treatment Effects and

### v3

@@ 3.2 Mechanisms and Outcomes
    $\widetilde{Y}=h(Y)$ where $h(\cdot)$ is a non-linear function;
then we call $\tilde{Y}$ an indirectly-affected outcome. Otherwise, $\widetilde{Y}$ is a directly-affected outcome.
The characterization of directly- and indirectly-affected outcomes is distinct from whether an outcome is or could be observed by the analyst. Consequently, our definition is distinct from the concepts of latent and observed outcomes (e.g., Fariss, Kenwick, and Reuning, 2020). In principle, directly-affected and indirectly-affected variables could both be latent and/or observed. We illustrate these possibilities concretely in Appendix D using three examples from (hypothetical) studies with some type of learning or updating mechanism. In empirical studies akin to our runnning example, the indirectly-affected outcome, vote choice (or self-reported vote choice), is generally observed whereas the directly-affected outcome, voter utility, is not. But this reflects a choice by researchers. Rather than measuring only vote choice, a researcher could invest in eliciting willingness-to-pay for the incumbent in an endline survey. This investment would render a measure of the directly-affected outcome observed.
The distinction between directly- and indirectly-affected outcomes represents a theoretical commitment about the causal sequencing of outcomes. In contrast, the distinction between latent and observed outcomes represents both a theoretical commitment about the relationship between outcomes and empirical concerns about what outcomes are or could be observed by the researcher. As we will highlight, the distinction directly- and indirectly-affected matters for the use of HTEs to provide evidence of mechanism activation. When outcomes later in a sequence are produced by a (non-zero) linear function of previous outcomes, they can be treated as directly-affected outcomes. This accomodates popular linear rescalings like $Z$-score transformations of observed outcomes.
The mediation framework that we build upon makes no distinctions based on the type or causal sequencing of outcomes. Yet, our motivating example reveals a distinction in what we can learn about mechanisms from HTEs on expected utility (the directly-affected outcome) versus vote choice (the indirectly affected outcome). This suggests that, in at least some cases, more structure is necessary to bridge the disjuncture between the mediation framework and many applied theories. This distinction between direct- and indirectly-affected outcomes is our solution to this disjuncture; there are certainly other ways to impose sufficient structure to bridge theoretical models and the statistical mediation framework. Our simple classification of outcomes proves quite useful for our analysis of what can be learned from HTEs.
4 HTEs and Mechanisms
When do heterogeneous treatment effects with respect to some covariate, $X_{k}$, provide evidence that a mechanism is active? To answer this question, we must first operationalize the concept of a theoretical mechanism. We view mediators—whether measured or unmeasured—as representations of a theoretical mechanism. For example, in the earlier example about exogenous shocks and voter behavior, the voter’s posterior belief represents the mediator. If a mediator is measured, researchers can, in theory, use off-the-shelf estimators to estimate the direct and indirect effects described above. Yet, there are two substantial limitations of mediation analysis. First, researchers may not have the ability to measure mediators. Not all mechanisms are measurable and even fewer are measured. Second, mediation analysis typically relies upon an additional ignorability assumption for identification. In our notation, this ignorability assumption, $Y_{i}(z^{\prime},m;X_{i})\perp M_{i}(z;X_{i})\mid Z_{i}=z;X_{i}$, holds that conditional on treatment and pre-treatment covariates $X_{i}$, potential outcomes are independent of the potential outcomes of the mediator (Imai, Keele, and Tingley, 2010). Critics allege that this assumption is unlikely to obtain and show that violation of this assumption generates bias in estimates of indirect effects (Gerber and Green, 2012). [^12] Strezhnev, Kelley, and Simmons (2021) relax sequential ignorability through an instrumental variables-based identification strategy and sensitivity analysis for assessing mediation effects.
The mediation framework allows us to precisely characterize HTEs with respect to covariates. However, it does not yet provide enough structure to link HTEs (or lack thereof) to mechanisms. To do so, we develop the concept of a mechanism detector variable (MDV) and impose an assumption. A MDV for a given mechanism induces a differential (average) causal effects through the mechanism of interest on the directly-affected outcome. This can be expressed in terms of (average) indirect effects. To economize notation, we will denote $AIE(X_{k}=x)=AIE(z,z^{\prime};X_{k}=x,X_{\neg k})$ as the average indirect effect of mechanism (mediator) when $X_{k}=x$.
Definition 4 (Mechanism detector variable (MDV)).
A pre-treatment covariate $X_{k}$ is a mechanism detector variable for mechanism $M$ if for some $x,x^{\prime}\in X_{k}$, $AIE(X_{k}=x)\neq AIE(X_{k}=x^{\prime})$ for the directly-affected outcome.
We then denote $\textbf{X}^{MDV}$ as the (possibly empty) set of all possible covariates that satisfy Definition 4 for the mechanism $M$ of a directly-affected outcome. Intuitively, if $X_{k}\in\textbf{X}^{MDV}$, then covariate $X_{k}$ can serve as an indicator for a mechanism/mediator of interest. [^13] A slightly stronger version of Definition 4 holds when $Y$ is continuously differentiable with respect to $M$ and $Z$ In this case, Definition 4 can be expressed as $\frac{\partial}{\partial X_{k}}\left(\frac{\partial Y}{\partial M}\frac{\partial M}{\partial Z}\right)\neq 0$. Under our definition of MDVs, it could be the case that $X_{k}$ moderates the effect of the treatment on the mediator. Interestingly, it could also be the case that $X_{k}$ moderates the effect of the mediator on the outcome. Both possibilities are depicted in Figure 2. In our running example, a variable measuring voters’ prior beliefs would be a MDV for the learning mechanism whereas a variable measuring valence assessments would not. This follows immediately from the DAG in Figure 1.
While theory—as in our running example—aids researchers in postulating MDVs for a given mechanism, empirical researchers seek to understand whether that mechanism is indeed active. Researchers then seek to detect evidence of a mechanism using treatment-by-covariate interactions with a proposed MDV. In other words, given a proposed MDV $X_{k}$, we want to learn whether $X_{k}\in\textbf{X}^{MDV}$.

### v4

@@ 3.2 Case 2: HTEs do not exist for outcome $Y(Z)$
    No MDV exists for mechanism $M_{1}$.
Proposition 3 shows that a lack of HTEs provides less information with regard to mechanism activation than is generally asserted. Under the exclusion assumptions, there are two reasons why HTEs may not exist with respect to a covariate, $X_{k}$. First, it may be the case that $X_{k}$ is not a MDV for mechanism $M_{1}$. In this sense, we have misspecified the theoretical relationship between a given covariate and a mechanism. Second, it may be the case that no MDV exists for mechanism $M_{1}$. As we discuss in Corollary 1, there are two possible reasons why a MDV would not exist for mechanism $M_{1}$. Importantly, we show that this could happen with an active or an inert mechanism $M_{1}$.
Corollary 1.
If no MDV exists for a mechanism $M_{1}$, there are two possibilities:
(1) Mechanism $M_{1}$ is not active.
(2) Mechanism $M_{1}$ is active, but produces the same effect for all units so there exists no $X_{k}$ for which $AIE^{Y}_{1}(X_{k}=x)\neq AIE^{Y}_{1}(X_{k}=x^{\prime})$.
Case (1) of Corollary 1 is implied by the definition of MDV. If a mechanism is inert—thereby producing an indirect effect of zero for all units—there cannot exist any MDVs, measured or unmeasured. In contrast, in Case (2), a mechanism can be active but it produces the same indirect effect for all units. Recall that a non-zero difference in average indirect effects at different levels of a covariate $X_{k}$ is a sufficient condition for mechanism activation. However, it is not a necessary condition for mechanism activation. These results show that, in contrast to standard interpretation, a lack of heterogeneity cannot tell us about whether a mechanism is active. Moreover, our theory could be misspecified, meaning that our postulated MDV, $X_{k}$ is not actually a MDV. An assessment of HTEs with respect to a single moderator cannot distinguish between these three possibilities. Nor can we assign probabilities to these (non-mutually exclusive) explanations without stronger assumptions.
3.3 Case 3: HTEs exist for transformed outcome $\widetilde{Y}(Z)$
To understand what HTEs reveal with respect to a (non-linearly) transformed outcome $\widetilde{Y}(Z)$, it is useful to introduce one final concept. We will denote $\textbf{X}^{R}=\{X_{1},X_{2},...,X_{K}\}$, as the set of all possible pre-treatment covariates with non-zero effects on the outcome, YY. [^16] Formally, if $X_{k}\in\textbf{X}^{R}$, then there exist $x\neq x^{\prime}\in X_{k}$ such that $Y(Z,M(Z;X_{k}=x,X_{\neg k});X_{k}=x,X_{\neg k})\neq Y(Z,M(Z;X_{k}=x^{\prime},X_{\neg k});X_{k}=x^{\prime},X_{\neg k})$. Covariates in $\textbf{X}^{R}$ can be thought of as “relevant” for predicting outcome $Y$. It is also useful to let X be the set of all possible pre-treatment covariates. It is clear that for any outcome, $Y$, and mechanism $M$, $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}\subseteq\textbf{X}$. Typically, these subsets will be proper.
We now return to our main question of interest: what do HTEs reveal with regard to mechanisms? Proposition 4 considers the case when there are HTEs in a covariate $X_{k}$. Here, we can learn that $X_{k}\in\textbf{X}^{R}$, but this is not informative about whether $X\in\textbf{X}^{MDV}$, since $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}$. Why do we observe heterogeneity in covariates in $\textbf{X}^{R}$ that are not MDVs? The non-linear transformation $h(\cdot)$ generically means that outcome $\widetilde{Y}$ does not satisfy Assumptions 1 and 2, even if $Y$ satisfies both assumptions. In this case, the difference in CATEs no longer identifies a difference in the conditional AIEs of interest! Indeed, the difference in CATEs for the transformed outcome is produced by the (sum of) differences in conditional $ADE^{\widetilde{Y}}$’s and conditional $AIE^{\widetilde{Y}}$’s for all mechanisms. Violation of the identifying assumptions means that we can no longer attribute HTEs to the mechanism of interest, $M_{1}$. This means that (absent stronger assumptions), we cannot learn about the activation of $M_{1}$ from HTEs on the transformed outcome.
Proposition 4.
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. Let observed outcome $\widetilde{Y}$ be a transformed outcome relative to $Y$. If HTEs exist with respect to $X_{k}$ for outcome $\widetilde{Y}$, then $X_{k}\in\textbf{X}^{R}.$
3.3.1 Case 4: HTEs do not exist for transformed outcome $\widetilde{Y}(Z)$
We now return to a final case of our theoretical analysis by asking when we are examining a transformed outcome that may be affected by mechanism $M_{1}$, what can we learn from a lack of HTEs? Proposition 5 indicates that in this case, we can infer that $X_{k}\in\textbf{X}$. This is obviously a vacuous result. We already know that $X_{k}\in\textbf{X}$ since $X_{k}$ is a covariate and X is the set of all covariates. However, the proposition tells us that we cannot assert that the absence of HTEs implies that $X_{k}$ is not a MDV. The logic here resembles the previous case. Non-linear transformation $h(\cdot)$ violates Assumptions 1 and 2, which means that we cannot attribute a lack of heterogeneity in CATEs to a lack of differences in a given conditional AIEs. We purposely state a vacuous result to emphasize how little can be ascertained about mechanisms from the lack of HTEs when outcomes are transformed non-linearly.
Proposition 5.
Suppose Assumptions 1-2 hold with respect to $X_{k}$ for outcome $Y$. Let observed outcome $\widetilde{Y}$ be a transformed outcome relative to $Y$. If HTEs do not exist with respect to $X_{k}$ for outcome $\widetilde{Y}$, then $X_{k}\in\textbf{X}.$
Often we make assumptions about the mapping $h$. For example, the mapping in (13) imposes assumptions about how latent attitudes translate into Likert-scale responses. When we are willing to make such assumptions, we can refine Proposition 5 slightly. Specifically, in Proposition 11, we show that if Assumptions 1 and 2 hold for an absolutely continuous $Y$, which is transformed by (13) (for any $Q\geq 2$ categories), if HTEs do not exist with respect to $X_{k}$, then $X_{k}\notin\textbf{X}^{R}$. Because $\textbf{X}^{MDV}\subseteq\textbf{X}^{R}$ we know then that $X_{k}\notin\textbf{X}^{MDV}$ if $X_{k}\notin\textbf{X}^{R}$. But as in Proposition 3 and Corollary 1, there are multiple possible explanations: our theory about how $X_{k}$ relates to mechanism $M_{1}$ could be wrong or no MDV exists for mechanism $M_{1}$. These possibilities mean that we cannot make an inference about mechanism (non)-activation from the absence of HTEs with respect to $X_{k}$.
3.4 Summary of results
In sum, our propositions characterize four cases into which we can classify attempts to ascertain mechanism activation from HTEs, as described in Table 2. This table suggests that in addition to the exclusion assumptions needed for identification of the difference of conditional average indirect effects of a mechanism (Assumptions 1 and 2), HTEs serve as a useful test of the activation of a mechanism when evaluated for an outcome that satisfies these exclusion assumptions. However even with these outcomes, we can make an inference mechanism activation in the presence of heterogeneity, but not in the absence of heterogeneity. In the next section, we consider how these results should inform applied theories and research design.
Case Outcome type HTEs? What can be inferred about mechanism $M_{1}$?
1 $Y$ Yes Difference in conditional indirect effect of mechanism $M_{1}$ is not zero $\Rightarrow$ Mechanism $M_{1}$ is active for at least one unit.
2 $Y$ No Three possible explanations (not mutually exclusive):
 (a) Covariate $X_{k}$ does not moderate the effect of mechanism $M_{1}$, regardless of whether the mechanism is active or not.
 (b) Mechanism $M_{1}$ is active and produces the same effect for all units. (No covariate moderates its indirect effect.)
 (c) Mechanism $M_{1}$ is inert (produces no effect) for all units.
$\Rightarrow$ No information about activation of mechanism $M$ without further assumptions.
3 $\widetilde{Y}$ Yes Covariate $X_{k}$ predicts outcome $\widetilde{Y}\Rightarrow$ No information about the activation of mechanism $M_{1}$.
4 $\widetilde{Y}$ No Without further assumptions, provides no information about relationship between $X_{k}$ and $\widetilde{Y}$. No information about activation of mechanism $M_{1}$.
Table 2: Summary of the interpretation of results. All cases assume that Assumptions 1 and 2 hold for covariate $X_{k}$, mechanism $M$, and outcome $Y$. Outcome $\widetilde{Y}$ is a transformed outcome relative to $Y$.
4 Applications and Guidance for Research Design
How should these results guide future efforts to test mechanisms using heterogeneous treatment effects? In this section, we use a set of four published applications to illustrate how these considerations should inform interpretation and prospective research design. In contrast to many methodological papers, our focus is on how theorized mechanisms underlying treatment effects link to a specific estimand: a difference in CATEs. [^17] These discussions generalize to a difference in conditional local average treatments (LATEs) or difference in conditional average treatment effects on the treated (CATTs). There are multiple ways to estimate this difference. For a randomized binary treatment, $Z_{i}$ and binary moderator $X_{ik}$—a candidate MDV—the most straightforward estimator of this difference in CATEs is $\beta_{3}$ in the OLS regression equation:
$Y_{i}$ $=\beta_{0}+\beta_{1}Z_{i}+\beta_{2}X_{ik}+\beta_{3}Z_{i}X_{ik}+\varepsilon_{i}$
For guidance on estimation, we refer readers to excellent treatments of estimation of HTEs (or interaction effects) including Brambor et al. (2006); Berry et al. (2009); Hainmueller et al. (2018). Additionally, see expositions of related inferential problems stemming from limited statistical power of interaction effects (e.g., McClelland and Judd, 1993) and the threat of multiple comparisons when HTEs in multiple moderators are estimated (e.g., Gerber and Green, 2012; Lee and Shaikh, 2014; Fink et al., 2014).
In line with our motivating example, we select the four applications that examine empirically partisan alignment (or bias) as a possible moderator for the effect of corruption information about an incumbent candidate. Appendix G.1 describes the contexts and designs of these studies. Figure 3 provides a summary of the theories invoked in each study. Panel (a)—which is analogous to our running model (while omitting valence)—suggests that information about corruption gives way to an analogous distaste for corruption mechanism. But voters also value the partisanship of the incumbent. Distaste for corruption (the mechanism) and partisanship enter voter preferences for the incumbent as distinct and additively separable considerations of voters. Panel (b) is equivalent to Panel (a), but includes a second mechanism linking corruption information to voter utility: a motivation of voters to coordinate votes within their precinct. Panel (c) returns to a single mechanism, voter distaste for corruption, but suggests that partisan alignment also conditions how voters process corruption information, suggesting the presence of motivated reasoning. Finally, Panel (d) includes only voter distaste for corruption and suggests that the degree to which voters are averse to corruption is a function of their partisan leanings. We justify our representation of each cited theoretical account in Appendix G.2. We note that the three papers with a single mechanism largely focus on voter distaste for observed corruption and emphasize the role of partisanship in moderating this effect. In contrast, Arias et al. (2019) focus on the second mechanism—voter coordination—whereas we will focus (largely) on the distaste mechanism in our analysis. We note that, unlike our model, these papers do not focus on corruption aversion as a moderator, though such variation is generally consistent with the theories they articulate (see Appendix G.2).